# AlabOS: A Python-based Reconfigurable Workflow Management Framework for Autonomous Laboratories

### v1

@@ Abstract
The recent advent of autonomous laboratories, coupled with algorithms for high-throughput screening and active learning, promises to accelerate materials discovery and innovation. As these autonomous systems grow in complexity, the demand for robust and efficient workflow management software becomes increasingly critical. In this paper, we introduce AlabOS, a general-purpose software framework for orchestrating experiments and managing resources, with an emphasis on automated laboratories for materials synthesis and characterization. We demonstrate the implementation of AlabOS in a prototype autonomous materials laboratory. AlabOS features a reconfigurable experiment workflow model, enabling the simultaneous execution of varied workflows composed of modular tasks. Therefore, AlabOS is well-suited to handle the rapidly changing experimental protocols defining the progress of self-driving laboratory development for materials research.
@@ 1 Introduction
Computational and data-driven approaches have shaped a new paradigm for materials discovery, leveraging recent advances in high-performance computing and machine learning. 1, 2, 3, 4, 5, 6, 7 Despite these advances, the experimental realization of computationally predicted compounds remains a slow, yet essential step in materials discovery and development8, 9, 10. Substantial progress has been made to increase experimental throughput while reducing human involvement, as demonstrated by several automated laboratories that utilize synthesis methods based on flow chemistry 11, 12, 13, 14, thin-film deposition 15, 16, solution-based synthesis 17, and solid-state synthesis 18, 19. These laboratories have also been integrated with predictive computational tools such as high-throughput density functional theory (DFT) calculations 20, 21, 1, 22, machine learning algorithms for automated interpretation of characterization data 23, 24, 25, 26, 27, and optimization algorithms that can perform decision making 28, 29, 30, 31, 32. The integration of computational modeling and artificial intelligence into autonomous laboratories gives rise to what is known as self-driving laboratories (SDLs), which can explore material spaces with minimal human intervention.33, 34, 35. The rising prominence of SDLs has elevated the significance of robotics and machine learning for of materials research. To accommodate this changing landscape from the human-centric experimentation diagram, improved tools are needed to manage robotics and the data they rapidly produce. 36, 37, 38
@@
In designing management software, one can learn from the automated workflows developed for high-throughput DFT calculations jain2015fireworks, 39, 40. These workflows provide a high-level user interface from which materials and calculation types can be specified, abstracting the lower-level tasks required to perform such calculations 41, 42. These computational management software need to effectively organize large datasets, a task similarly required for experimental management. However, orchestrating an autonomous laboratory creates additional challenges, in particular, requiring the seamless integration of computations and data with physical hardware and experimental samples, which is crucial for maintaining uninterrupted operation.
With the increasing prevalence of laboratory automation, a range of workflow management programs, such as ChemOS 2.0 43, ARChemist 44, HELAO 45 and more 46, 47, 48, have been developed for SDLs. These pioneering efforts are designed to tackle the lack of flexibility and extendability in experimental management in autonomous laboratories by providing a general framework to manage the experiment workflows under different hardware settings. They have greatly facilitated the management of workflows in autonomous laboratories with minimal programming skill requirements for researchers to build an autonomous laboratory. However, these frameworks tend to focus on a fixed experimental workflow (e.g., sequential deposition of samples onto a thin film substrate, followed by characterization of those samples) and can face challenges in accommodating workflows with more complex ordering of samples and operations that involve parallelization (e.g., annealing multiple samples in one furnace) and tree-like dependencies (e.g., parallel XRD and SEM/EDS analysis after synthesis).
@@
Herein, we present AlabOS, a versatile and accessible workflow management framework for autonomous laboratories. Inspired by the Robot Operating System (ROS) 49, this system provides high-level Application Programming Interfaces (APIs) to define devices in autonomous laboratories and create customized workflows with a high degree of flexibility. This Python-based software is platform-independent and designed to be user-friendly, requiring only minimal programming skills for community adoption. An experimental workflow is comprised of multiple modular tasks, which are organized in a directed graph to represent the dependencies between tasks. To eliminate possible conflicts between concurrent tasks, a resource management solution assigns lab resources to certain tasks when needed. Considering the errors and maintenance demands in the daily operations of an autonomous laboratory, a status monitoring and notification system is built into the software along with a browser-based graphic user interface (GUI). AlabOS serves as a general framework for managing workflows, designed to simplify the programming required to establish an autonomous laboratory. At the same time, it maintains the flexibility to accommodate various workflows within the laboratory.
@@
The AlabOS system is actively deployed in the A-Lab 28, an autonomous laboratory for inorganic materials synthesis housed at Lawrence Berkeley National Laboratory (LBNL). At the time of writing, it has synthesized and characterized over 2,500 distinct samples under the control of AlabOS.
@@ 2.1 Status representations
A device refers to a piece of hardware to process or collect data from sample(s). Each device entity is linked to hardware in the lab that can send commands, perform physical operations, and collect data. Some examples of these operations in materials synthesis include sample dispensing, weighing, heating, and grinding. In AlabOS, a device is defined in Python code that provides the methods to communicate with the data transfer protocols like MODBUS 50, TCP/IP, and serial. Each device will have an entry in the database that indicates its status: whether it is occupied by a task, or whether it is paused for maintenance.
@@ 2.2 AlabOS architecture
To interface with the manager processes in a user-friendly way, a dashboard server (Fig. 2c) with a GUI is provided. This allows the operator to control and monitor the progress of any experiments running in the laboratory, while also showing the current states of all devices in the lab. In addition to the GUI, this dashboard provides an API that can receive new experiment submissions in a JSON format 51. Each experiment submission is validated using Pydantic 52 models to ensure the correctness of all formats and values.
@@ 3.1 Resource management
Unlike traditional materials research laboratories, where human experts perform various tasks (like sample preparation, heating, and characterization) in a serial fashion, autonomous laboratories usually take a more distributed and parallelized approach. For example, a robot may dispense and weigh precursor powders while an automated furnace is heating a separate batch of samples that have already been weighed. There is typically a queue of pending tasks that the laboratory must plan and account for. To ensure the lab’s continuous and successful operation without requiring human intervention, the system should avoid any possible conflicts between these tasks while also maximizing the overall throughput by organizing the distribution of tasks in a parallel fashion.
In AlabOS, resource management is carried out using a cooperative multitasking schema 53. Each task proactively reserves the resources needed to complete its operations. Once those operations are finished, the associated resources are released to be assigned to other tasks. To avoid the occurrence of “dead resources” where tasks are completed without releasing their resources properly, the system introduces a with-statement context for all requests, as shown in Fig. 3. Once resources are assigned to a task, they are recorded in a resource request context. All operations involving these resources need to be performed within this context to avoid a permission error. When the task is completed and exits the request context, it will automatically release all the requested resources, regardless of the task’s success or failure.
@@
A resource request can contain multiple devices and sample positions. In an autonomous laboratory, it is sometimes common to have multiple devices operating with the same functions (e.g., multiple furnaces for heating) to increase the lab’s throughput. A task can operate on equivalent devices to obtain the desired results. In the task definition, a request can either specify the name of an exact device to use (e.g., Furnace A) or a type of device (e.g., muffle furnace). If a device type is specified, the resource manager will find any available device(s) under that type. The sample position request is then attached to the specific device. For example, one task may want to reserve the sample positions inside one furnace chamber. In this case, the sample positions request will be handled once the device request can be satisfied. When the resource manager finds an available combination of devices that can satisfy the request, it will advance to solve the available sample positions based on the proposed devices. Once both the devices and sample positions requests can be met, a request will be fulfilled and assigned to the corresponding tasks.
@@
When initiating a resource request, a priority can also be attached to determine the urgency of the request. The priority is encoded as an integer ranging from 1 to 100. The default priority of all the tasks is 20, which indicates normal priority. When the resource manager polls the pending requests, it will first rank them by their priority and then by their submission time. The request that has a higher priority and was submitted earlier will generally be handled first.
@@ 3.2.2 Notification
Human interaction in an autonomous laboratory is essential to maintain a smooth operation. When something unexpected is detected in the laboratory (e.g., increased environment temperature, or robot arm failing to pick up a sample), the autonomous lab should notify the operators in time to avoid further damage. Clear instructions should be given so that it requires minimal knowledge to act according to the notification.
In AlabOS, the notification can be initiated by either the system (manager instances) or the task processes. It can be used whenever a human action is needed. Each notification contains a prompt field and a list of possible options for the operators to select. The notification requires a receipt of acknowledgment from the operators with one option selected. Once that option is selected and the operator has completed the necessary tasks, the program will be carried out according to the selected actions.
Apart from the dashboard messages, notifications can also be configured to send over Slack bot or email. Further extensions can be made to other notification platforms, such as IFTTT 54, with minor programming efforts.
3.2.3 Error handling
Because autonomous laboratories are generally used for research purposes, errors in hardware and software are inevitable. For example, the robot arm encountering an unexpected object can lead to a hardware error, while the driver code failing to communicate with the robot arm would be considered a software error. In AlabOS, errors are raised as standard Python exceptions in the task actor process. The errors are classified as recoverable or unrecoverable. Recoverable errors refer to issues caught in the task process, usually in Python’s “try-except” block. The raised error will not lead to the exiting of a program. Instead, there will be a handling routine that either retries or notifies human operators to check and recover manually. In contrast, an unrecoverable error will lead to the failure of an entire task. These are not caught by the “try-except” block. AlabOS provides a default handling routine to notify human operators of all unrecoverable errors. They can then remove and replace any affected samples, reset the occupied devices, and resume the system to complete any other remaining samples.
@@ 3.3 Device control
AlabOS system uses a centralized manager to monitor and control all devices in a laboratory. The task actors do not communicate directly with the physical equipment to execute commands and read data. Instead, they send all commands to the device manager through a remote procedure call (RPC) 55. In each task, an identical RPC proxy class will be created as an in-place replacement for the task to send commands. In this way, the RPC communication will not change the way driver methods are called in the task definition, as opposed to directly calling the devices’ methods.

### v2

@@ Abstract
The recent advent of autonomous laboratories, coupled with algorithms for high-throughput screening and active learning, promises to accelerate materials discovery and innovation. As these autonomous systems grow in complexity, the demand for robust and efficient workflow management software becomes increasingly critical. In this paper, we introduce AlabOS, a general-purpose software framework for orchestrating experiments and managing resources, with an emphasis on automated laboratories for materials synthesis and characterization. AlabOS features a reconfigurable experiment workflow model and a resource reservation mechanism, enabling the simultaneous execution of varied workflows composed of modular tasks while eliminating conflicts between tasks. To showcase its capability, we demonstrate the implementation of AlabOS in a prototype autonomous materials laboratory, A-Lab, with around 3,500 samples synthesized over 1.5 years.
@@ 1 Introduction
Computational and data-driven approaches have shaped a new paradigm for materials discovery, leveraging recent advances in high-performance computing and machine learning. 1, 2, 3, 4, 5, 6, 7 Despite these advances, the experimental realization of computationally predicted compounds remains a slow, yet essential step in materials discovery and development8, 9, 10. Substantial progress has been made to increase experimental throughput while reducing human involvement, as demonstrated by several automated laboratories that utilize synthesis methods based on flow chemistry 11, 12, 13, 14, thin-film deposition 15, 16, solution-based synthesis 17, and solid-state synthesis 18, 19. These laboratories have also been integrated with predictive computational tools such as high-throughput density functional theory (DFT) calculations 20, 21, 1, 22, machine learning algorithms for automated interpretation of characterization data 23, 24, 25, 26, 27, and optimization algorithms that can perform decision making 28, 29, 30, 31, 32. The integration of computational modeling and artificial intelligence into autonomous laboratories gives rise to what is known as self-driving laboratories (SDLs), which can explore material spaces with minimal human intervention33, 34, 35. The rising prominence of SDLs has elevated the significance of robotics and machine learning for materials research. To accommodate this changing landscape from the human-centric experimentation diagram, improved tools are needed to manage robotics and the data they rapidly produce. 36, 37, 38
@@
In designing management software, one can learn from the automated workflows developed for high-throughput DFT calculations 39, 40, 41. These workflows provide a high-level user interface from which materials and calculation types can be specified, abstracting the lower-level tasks required to perform such calculations 42, 43. These computational management software need to effectively organize large datasets, a task similarly required for experimental management. However, orchestrating an autonomous laboratory creates additional challenges, in particular, requiring the seamless integration of tasks and data with physical hardware and experimental samples, which is crucial for maintaining uninterrupted operation.
Several workflow management programs have previously been developed and implemented in different autonomous laboratories with diverse applications. For example, ChemOS 2.0 44 proposes an integral platform between DFT calculation, Bayesian optimization, and automated equipment communicated through SiLA2 protocol 45. Helao 46 and Helao-async 47 present a highly modularized design by implementing each component as a web server, enabling workflow management across laboratories. Bluesky 48 has been applied to synchrotron facilities, with a strong emphasis on controlling synchrotron characterization hardware, to improve data collection efficiency but not targeted to materials synthesis. Additional software packages have supported laboratories involving automated experiments, featured with expandability and reconfigurability across different experiment stations 49, standardized data specification and storage 50, and seamless integration with other science tools as a “science factory” 51. These software solutions have successfully orchestrated experiment workflows in autonomous laboratories within their domains.
On the other hand, recent trends toward increased throughput and greater complexity in autonomous laboratories necessitate workflow software to manage more diverse workflows with a higher volume of autonomous devices. For instance, in our use case as well as others aiming to boost overall throughput, a laboratory can feature multiple equivalent devices for a single task, such as several furnaces for heating or two X-ray diffractometers for characterization. Due to the lack of knowledge about the availability of devices during runtime, the workflow management system must dynamically allocate the devices to each task based on their needs, thereby maximizing the utilization rate of the autonomous laboratory. Moreover, as general-purpose research laboratories often run different workflows with varied task sequences simultaneously, a reconfigurable workflow model has become essential, with certain parts of the workflow being executed in parallel. For example, after a sample is synthesized, multiple characterizations can be performed concurrently while storage operation has to wait until all the characterizations are done. Depending on the needs of the experiments, different experiment procedures need to be executed in the same laboratory setting. Thus, an expressive workflow model is required to effectively encode such task order dependencies while preserving flexibility when composing the workflow. Efforts have been made to handle some of these challenges. For example, a scheduling strategy is developed to tackle the multi-device problem, which requires an accurate time model for each task to ensure efficiency.52 However, only by addressing all of these practical needs, a workflow management software can fit better into an autonomous laboratory with increased complexity and throughput and allow researchers greater freedom to push the boundaries of autonomous experimentation.
@@
Herein, we present AlabOS, a versatile and accessible workflow management framework for autonomous laboratories. The system features a graph-based experimental workflow model with tasks being the nodes and task dependencies being the edge. A central resource manager is built into the system, to track the status of devices and allocate resources properly to each task, thus eliminating possible conflicts between tasks running at the same time in the laboratory. Furthermore, AlabOS proposes the concept of sample position, representing a position in the laboratory that can hold one sample. By tracking the sample’s position, AlabOS makes it possible to track individual samples in real time. This Python-based software is platform-independent and designed to be user-friendly, requiring basic knowledge about database and parallel programming for community adoption. Considering the errors and maintenance demands in the daily operations of an autonomous laboratory, a status monitoring and notification system is built into the software along with a browser-based graphic user interface (GUI). AlabOS serves as a general framework for managing workflows, designed to simplify the programming required to establish an autonomous laboratory. At the same time, it maintains the flexibility to accommodate various workflows within the laboratory. Other features of AlabOS include:
  • Provide a solution for autonomous laboratory workflow management, with common functionalities like resource management, and notification built in.
  • Built with MongoDB (NoSQL) backend, supporting a flexible schema, and allowing standardized continuous development of each task input, output, and data storage operations.
  • A standard way to define devices and tasks with base classes when setting up a new lab with AlabOS. A simulation mode is also built in, allowing quick debugging of the devices and tasks.
  • Submission and status monitoring APIs with JSON format. The users can set up scripts for job submission and queries on top of AlabOS. The input and output of the experiments can be validated by Pydantic 53 before being stored in the database.
@@
The AlabOS system is actively deployed in the A-Lab 18, an autonomous laboratory for inorganic materials synthesis housed at Lawrence Berkeley National Laboratory (LBNL). At the time of writing, it has synthesized and characterized over 3,500 distinct samples under the control of AlabOS.
@@ 2.1 Status representations
A device refers to a piece of hardware to process or collect data from sample(s). Each device entity is linked to hardware in the lab that can send commands, perform physical operations, and collect data. Some examples of these operations in materials synthesis include sample dispensing, weighing, heating, and grinding. In AlabOS, a device is defined in Python code that provides the methods to communicate with the data transfer protocols like MODBUS 54, TCP/IP, and serial. Each device will have an entry in the database that indicates its status: whether it is occupied by a task, or whether it is paused for maintenance.
@@ 2.2 AlabOS architecture
To interface with the manager processes in a user-friendly way, a dashboard server (Fig. 2c) with a GUI is provided. This allows the operator to control and monitor the progress of any experiments running in the laboratory, while also showing the current states of all devices in the lab. In addition to the GUI, this dashboard provides an API that can receive new experiment submissions in a JSON format 55. Each experiment submission is validated using Pydantic 53 models to ensure the correctness of all formats and values.
@@ 3.1 Resource management
Unlike traditional materials research laboratories, where human experts perform various tasks (like sample preparation, heating, and characterization) serially, autonomous laboratories usually take a more distributed and parallelized approach. For example, a robot may dispense and weigh precursor powders while an automated furnace is heating a separate batch of samples that have already been weighed. There is typically a queue of pending tasks that the laboratory must plan and account for. To ensure the lab’s continuous and successful operation without requiring human intervention, the system should avoid any possible conflicts between these tasks while also maximizing the overall throughput by organizing the distribution of tasks in a parallel fashion.
In AlabOS, resource management is carried out using a cooperative multitasking schema 56. Before tasks perform any operations in the laboratory, it has to request the necessary resources from AlabOS system. All the resource requests are handled by a module named resource manager, which will parse each request and check if there are any idle devices and sample positions that can fulfill the request. Once the request can be fulfilled, the resource manager will mark the assigned devices and sample positions as occupied and let the task know which devices and sample positions it can use. The task can then send commands to the assigned devices and move samples into the assigned sample positions. Once those operations are finished, the associated resources are released to be assigned to other tasks. To avoid the occurrence of “dead resources” where tasks are completed without releasing their resources properly, the system introduces a with-statement context for all requests, as shown in Fig. 3. Once resources are assigned to a task, they are recorded in a resource request context. All operations involving these resources need to be performed within this context to avoid a permission error. When the task is completed and exits the request context, it will automatically release all the requested resources, regardless of the task’s success or failure.
@@
A resource request can contain multiple devices and sample positions. In an autonomous laboratory, it is sometimes common to have multiple devices operating with the same functions (e.g., multiple furnaces for heating) to increase the lab’s throughput. A task can operate on equivalent devices to obtain the desired results. In the task definition, a request can either specify the name of an exact device to use (e.g., Furnace A) or a type of device (e.g., muffle furnace). If a device type is specified, the resource manager will find any available device(s) under that type. The sample position request is then attached to the specific device. For example, one task may want to reserve the sample positions inside one furnace chamber. In this case, the sample positions request will be handled once the device request can be satisfied. When the resource manager finds an available combination of devices that can satisfy the request, it will advance to solve the available sample positions based on the proposed devices. Once both the devices and sample positions requests can be met, a request will be fulfilled and assigned to the corresponding tasks. With such a non-preemptive resource allocation strategy, the resource request serves in a “first-come-first-serve” strategy with the task priority taken into consideration. This does not need additional information like how long the task will occupy these resources. However, it also assumes that all the tasks do not occupy resources for an unreasonably long time which would lead to sample pileups in the lab. It is generally recommended for the laboratory developer to have a good knowledge of the bottlenecks and throughput of the laboratory and set up the right amount of tools to avoid this. On the other hand, it is always a good idea to break the resource request into many small pieces to avoid booking some resources for too long a time. Before the task definition is deployed in the physical laboratory, the lab developer should fully test the workflow definition in the simulation mode described in Section 3.3 to ensure the expected behavior of a task definition. More sophisticated scheduling of equipment reservation techniques 52 could be integrated within AlabOS to further optimize the throughput.
@@
When initiating a resource request, a priority can also be attached to determine the urgency of the request. The resource request’s priority is kept the same as the priority of the task, which is specified by the user during submission as a task parameter (default to “NORMAL” priority). The resource requests’ priority can also be overridden by supplying a customized value when creating a resource request inside the task. It is sometimes useful to force a higher priority to the post-processing operations in a task. For example, the sample must be unloaded from the X-ray diffractometer before the next sample can be loaded. In this case, it is necessary to prioritize the unload operation to avoid gridlock in the laboratory. The priority is encoded as an integer ranging from 1 to 100. The default priority of all the tasks is 20, which indicates normal priority. When the resource manager polls the pending requests, it will first rank them by their priority and then by their submission time. The request that has a higher priority and was submitted earlier will generally be handled first.
In an autonomous laboratory, many tasks can request resources at the same time. It is necessary to handle the large amount of possible resources on time so that the task can be executed faster. To understand the performance of the resource manager, a virtual lab with $M$ devices, each of which has 20 associated sample positions, is defined. At each step, there are $N$ tasks that request one device with one to twenty sample positions. Such configuration simulates the situation when the laboratory has a heavy workload. At each step, the resource manager will try to assign resources to each task. The CPU wall time is collected at different conditions. As shown in Figure. 4(a), when the number of devices is fixed, the processing time for all the tasks’ requests scales linearly with the task number at small task numbers ($\leq$ 120). The average processing time for one task is 7.407 ms obtained from the slope of the fitted line. When the number of tasks continues to increase, the processing time deviates down from the fitted line. This may be attributed to a saturation of the resources. When there are way more tasks requesting resources compared to the available devices, most tasks’ requests cannot be fulfilled in one step, leading to a processing time shorter than the time predicted by the linear relationship. The performance of the resource manager on different numbers of devices is shown in Figure. 4(b). Similar to the performance under varied task numbers, with the number of devices increasing, the processing time first follows a linear relationship with a slope of 14.073 ms/device and then becomes flat due to the saturation of the tasks. When the devices are much more than the tasks, most devices remain idle and do not cost any time for resource assignment.
Compared to the time required for completing a normal operation in the laboratory for solid-state synthesis, which usually takes several minutes to hours, the time for assigning resources is nearly negligible, thus ensuring a higher turnover rate in the laboratory. In some cases, multiple operations can be short and complete in a few seconds, while requiring different resources to complete. If that is the case, it is generally recommended to combine these short operations and request the resources all at once to minimize the waiting time.
Figure 4: The execution time of processing resource requests under different numbers of tasks, devices, and sample positions. The time to process all the resource requests (a) with varied numbers of tasks in a virtual lab with 50 devices and 1000 sample positions (20 sample positions per device), (b) with fixed 200 tasks in a virtual lab with varied number of devices and sample positions (20 sample positions per device). The test is run on a Macbook Pro 2021 with M1 Pro CPU, 32 GB RAM. Each point is run 10 times repeatedly to obtain the error bar. The top and bottom of the error bar represent the 25th and 75th quantile of all the runtime data.
@@ 3.2.2 Notification
Human interaction in an autonomous laboratory is essential to maintain a smooth operation. When something unexpected is detected in the laboratory (e.g., increased environment temperature, or the robot arm failing to pick up a sample), the autonomous lab should notify the operators in time to avoid further damage. Clear instructions should be given so that it requires minimal knowledge to act according to the notification.
In AlabOS, the notification can be initiated by either the system (manager instances) or the task processes. It can be used whenever a human action is needed. Each notification contains a prompt field and a list of possible options for the operators to select. The notification requires a receipt of acknowledgment from the operators with one option selected. Once that option is selected and the operator has completed the necessary tasks, the program will be carried out according to the selected actions. With clear notification messages designed when defining the tasks, the notification can be handled by the users without much knowledge about the autonomous laboratory. Thus, the maintenance jobs can be distributed among the lab developers and lab users.
Apart from the dashboard messages, notifications can also be configured to send over Slack bot or email. Further extensions can be made to other notification platforms, such as IFTTT 57, with minor programming efforts.
3.2.3 Exception handling
Because autonomous laboratories are generally used for research purposes, exceptions in hardware and software are inevitable. For example, the robot arm encountering an unexpected object can lead to a hardware exception, while the driver code failing to communicate with the robot arm would be considered a software exception. In AlabOS, errors are raised as standard Python exceptions in the task actor process. The exceptions are classified as recoverable or unrecoverable. Recoverable exceptions refer to issues caught in the task process, usually in Python’s “try-except” block. The raised error will not lead to the exiting of a program. Instead, there will be a handling routine that either retries or notifies human operators to check and recover manually. In contrast, an unrecoverable exception will lead to the failure of an entire task. These are not caught by the “try-except” block. AlabOS provides a default handling routine to notify human operators of all unrecoverable exceptions. They can then remove and replace any affected samples, reset the occupied devices, and resume the system to complete any other remaining samples.
@@ 3.3 Device control
AlabOS system uses a centralized manager to monitor and control all devices in a laboratory. The task actors do not communicate directly with the physical equipment to execute commands and read data. Instead, they send all commands to the device manager through a remote procedure call (RPC) 58. In each task, an identical RPC proxy class will be created as an in-place replacement for the task to send commands. In this way, the RPC communication will not change the way driver methods are called in the task definition, as opposed to directly calling the devices’ methods.